# The impact of the redshift-dependent selection effect of halos on the redshift-space power spectrum

### v1

@@ 1 Introduction
A standard approach to galaxy clustering analysis involves the use of the two-point correlation function or its Fourier-transformed counterpart, the power spectrum. The baryonic acoustic oscillations (BAO) [9] and the apparent anisotropic clustering due to the use of an incorrect cosmological model, i.e. the Alcock-Paczynski (AP) effect [10], in the measured two-point statistics provide a powerful geometrical probe of cosmological distances [11, 12]. In addition, peculiar velocities of galaxies cause characteristic anisotropies in redshift-space distribution of galaxies, i.e. redshift-space distortion (RSD) effect [13], and can be used to probe the growth of structure, free of galaxy bias uncertainty [14, 15]. More recently, advancements in the theory and methods of galaxy clustering analysis have enabled a more comprehensive approach, known as the “full-shape” analysis [16, 17]. This method incorporates BAO, AP, and RSD information as well as the shape information of the underlying matter power spectrum up to the quasi nonlinear scales within an assumed theoretical framework, such as the $\Lambda$CDM model.
@@
Therefore, the purpose of this paper is to investigate the impact of the selection effect on the redshift-space power spectrum. To achieve this, in this paper, we consider a case where an assumed sample of galaxies follows a redshift distribution, denoted as $n(z)$ or $n(\chi)$, representing the number density of galaxies per unit redshift interval or per unit radial comoving distance interval. We do not consider the redshift evolution of galaxy clustering (structure formation) to focus on the impact of the redshift-dependent selection effect. Hence in our setting, the constant comoving number density, $n(\chi)\mathrm{d}V(\chi)={\rm const.}$, corresponds to a homogeneous sample or the absence of the selection effect. We first use the linear theory to show that the selection function generally causes a bias in the multipole moments of the redshift-space galaxy power spectrum, along the Feldman-Kaiser-Peacock estimator [21]. We then use halo catalogs from cosmological N-body simulations, AbacusSumitt [22, 23], to construct mock catalogs of galaxies (halos), by selecting halos based on a redshift-dependent mass threshold such that the resulting redshift distribution of the selected halos follows the target $n(z)$ for the SDSS BOSS LOWZ- or CMASS-like galaxies. We measure the multipole moments of the redshift-space power spectrum from the mock catalogs and then compare the moments with those obtained from halo catalogs of a single mass threshold, i.e. single tracers, in the same simulations. This comparison quantifies the impact of the selection effect on the redshift-space power spectrum.
@@ 2.1 N-body Simulations and Halo Catalogs
We use the halo catalogs constructed from N-body simulation data for the Planck 2018 $\Lambda$CDM model, provided by AbacusSummit [22, 23].  [^2] https://abacussummit.readthedocs.io/en/latest/citation.html The N-body simulations were performed using $6912^{3}$ particles in a comoving cubic box with side length of $2h^{-1}\mathrm{Gpc}$. The particle mass was $2\times 10^{9}h^{-1}\mathrm{M_{\odot}}$ and the force softening scale was $7.2h^{-1}\mathrm{kpc}$. In this paper we use only the halo catalog and do not use the N-body particle data. In particular, we use the halo catalogs at a single redshift output, $z=0.5$, because we focus on the impact of selection effect on clustering measurements and, in other words, we do not include the effect of redshift evolution in order to highlight our results. Halos in each simulation realization were identified using CompaSO halo finder [24]. In this paper we use only central halos (labeled “L0” in the released catalogs). We use, as the position of each halo, the center-of-mass of the member particles and, as the velocity of each halo, the center-of-mass velocity for each halo, as provided in the halo catalogs. We use 25 realizations of the halo catalogs to reduce the statistical errors in our results.
@@ 2.2.1 Mock catalogs with galaxy selection effect
To mimic actual galaxy surveys, we consider the LOWZ and CMASS galaxy samples of Sloan Digital Sky Survey Baryon Oscillation Spectroscopic Survey (BOSS) Data Release 11 (DR11) [^3] https://data.sdss.org/sas/dr11/boss/lss/; we consider LOWZ galaxies at $z>0.09$, and CMASS galaxies at $z>0.40$, respectively, as shown in Fig. 1. In addition to these cases we also consider a selection function of galaxies that follows a power-law selection of $n(z)$, given by $n(z)\propto z^{-\alpha}$, for the generality of our discussion. We investigate three cases, $\alpha=0.5,1.0,2.0$, and determine their normalizations so that the resulting averaged number density in the radial distance range of $[0,1000]h^{-1}$pc, which roughly corresponds to the range of LOWZ, becomes the same as in the LOWZ case (see the right panel in Fig. 1). Throughout this paper we assume a distant observer approximation, and take the $z$-axis of each simulation box to be along the line-of-sight direction.
@@
Figure 1: The number density of LOWZ (left panel) and CMASS (middle panel) galaxies as a function of redshift, denoted as $n(z)$, taken from the SDSS DR11 catalog. Note that $n(z)$ is the number density per unit redshift interval, and we here plot $n(z)$ as a function of the comoving radial distance using the inverse relation of redshift and the comoving radial distance, $z=z^{-1}(\chi)$, for Planck 2018 cosmology. The LOWZ galaxies are taken from the redshift range $z\simeq[0.09,0.5]$, while the CMASS galaxies are from the range $z\simeq[0.4,0.8]$. For the sake of illustration, we set the comoving radial distance at the lowest redshift to $\chi=0$ for each galaxy sample. We also consider a case of $n(z)$ following a power-law scaling given as $n(z)\propto z^{-\alpha}$ (right panel) to study the impact of galaxy selection effect as a general case.
@@
The LOWZ and CMASS galaxies are selected based on the color and magnitude cuts [19, 25, 26, 20]. It is very difficult and nearly impossible to exactly mimic the actual color-magnitude cut for the galaxy selection using the halo catalog, because it requires simulating galaxies while accounting for the complex physics of galaxy formation and evolution. In this paper, we employ the following simplified selection function as a working example. For a given $n(z)$ for the sample of galaxies considered, we select halos with masses above the mass threshold $M_{\rm th}(z)$ in each simulation realization so that the resulting number density matches the $n(z)$ in each redshift bin $[\chi,\chi+\mathrm{d}\chi]$ (since the comoving radial distance is given by $\chi=\chi(z)$). We determine the mass threshold $M_{\rm th}(z)$ in each redshift bin from the average of the mass functions from 25 realizations of halo catalog, to minimize the sample variance effect. Since heavier halos tend to host more massive and therefore brighter galaxies [27], one could consider that this mass-threshold selection at each redshift mimics to some extent a flux limited sample of galaxies. In this paper, we refer to this method as abundance matching (AM) method, and denote the power spectrum measured from this catalog as $P_{\rm AM}(k)$. Fig. 2 shows the mass thresholds as a function of redshift, $M_{\rm th}(z)$, which we use to construct mock catalogs for LOWZ, CMASS and power-law samples, respectively.
@@
The mock galaxy catalogs are distributed in a rectangular-shape region of $2^{2}\times 1~{}(h^{-1}{\rm Gpc})^{3}$ volume. We generate two mocks from the halo catalog in each N-body simulation of $8~{}(h^{-1}{\rm Gpc})^{3}$ volume. We apply zero padding to the halo distribution to perform the Fast Fourier Transform in a $2^{3}~{}(h^{-1}{\rm Gpc})^{3}$ cubic box and then estimate the power spectrum using the Feldman-Kaiser-Peacock (FKP) estimator as described in Section 2.3 in detail. Since the mock catalog violates the periodic boundary condition, the estimated power spectrum is affected by the window convolution. For the following results, we use 50 realizations of the mock catalogs for each galaxy sample.
@@ 2.2.2 Mock catalogs with a single halo-mass threshold
To quantify the galaxy selection effect, we make a mock catalog which has no galaxy selection effect compared to the AM catalog. We create such a galaxy catalog as follows:
@@
    We construct different halo catalogs as a function of varying mass thresholds $M_{\rm th}$ in each realization.
@@
    For each halo catalog of a given mass threshold, we randomly select halos in each redshift bin (or along the $z$-axis direction in our setting) so that the resulting redshift distribution of the number density follows $n(z)\propto\bar{n}_{\rm g}(z)$, where $\bar{n}_{\rm g}(z)$ is the redshift distribution for the LOWZ, CMASS or power-law galaxy sample, respectively, as shown in Fig. 1. The power spectrum estimated using the FKP estimator (see below) retains the original power spectrum of the single mass threshold, and also incorporates the similar survey window effect as in the $P_{\rm AM}(k)$ for each of the three galaxy samples.
  • (iii)
    We find the halo catalog of a certain mass threshold such that its (real space) power spectrum matches the target spectrum $P_{\rm AM}(k)$ for each of the three galaxy samples on the linear scales in the range $k=[0.02,0.1]~{}h{\rm Mpc}^{-1}$. In this step we perform the $\chi^{2}$ fitting taking into account the sample variance error in each $k$ bin estimated from the variance of the 50 realizations.
In the following we call this halo catalog as a single mass-threshold sample that is free from the selection effect. We will compare the two power spectra measured from the AM sample and the single mass-threshold sample to quantify the impact of the galaxy selection effect.
Orange line in each panel of Fig. 3 shows $n(z)$ of a single mass-threshold sample, constructed based on the above method, for the LOWZ or CMASS galaxy sample. The dashed line in each panel shows the number density of halos corresponding to a certain mass threshold, which by definition has a constant amplitude in the $z$ direction because we are using the halo catalogs at the single redshift output ($z=0.5$). Then, we randomly select halos in each radial distance bin to satisfy $n(z)\propto n_{\rm LOWZ}(z)$ or $n_{\rm CMASS}(z)$, respectively. The proportional factor is determined by the average of all the redshift bins. The power spectrum measured from this single mass-threshold sample includes the survey window effect similarly to that for the AM galaxy sample (see Section 2.3). Note that the shot noise contamination estimated by the FKP estimator is different between the AM and single mass-threshold samples, and we subtract it from the measured power spectrum. The single mass-threshold sample in Fig. 3 has the power spectrum with the closest amplitude to the spectrum from the AM sample on the linear scales (see the above step (iii)), for each of the galaxy samples. For LOWZ sample, the single mass threshold is $M_{\rm th}=9.14\times 10^{12}h^{-1}{\rm M_{\odot}}$, and the proportional factor in $n(z)\propto n_{\rm g}(z)$ is about $1/4$. For CMASS sample, the single mass-threshold is $M_{\rm th}=1.39\times 10^{12}h^{-1}{\rm M_{\odot}}$, and the proportional factor is about $1/2$.
@@
Figure 3: A figure illustrating the method for incorporating the window effect on the single-mass threshold sample, used to quantify the impact of selection effects in comparison to the results of the AM catalog. For the single mass-threshold sample, we randomly select halos above the single mass threshold in each redshift bin such that the redshift dependence of the number density is proportional to that for the LOWZ, CMASS or power-law sample in Fig. 1 (here we show results only the for the LOWZ and CMASS samples) (see Section 2.2.2 for details on the single mass-threshold sample). This single mass-threshold hold sample has the same window function as that for the AM sample.
@@ 2.3 FKP estimator
where $\bar{n}(\mathbf{x})$ is given by the redshift distribution $\bar{n}_{\rm g}(z)$ (Fig. 1) (our mocks do not have any variations in the mean number density in the $x,y$-directions), and we adopt $P_{0}=10^{4}~{}h^{-3}\rm{Mpc}^{3}$ for the AM sample. For the single mass-threshold sample, we set $P_{0}=10^{4}/\beta~{}h^{-3}\rm{Mpc}^{3}$ in Eq. (2.2), where the $\beta$ is the proportional factor of $\bar{n}_{M_{\rm th}}(z)\propto\bar{n}_{\rm g}(z)$ in Section 2.2.2.
@@
$\braket{\hat{P}(\mathbf{k})}$ =|$=\braket{}{F(\mathbf{k})}{{}^{2}}-P_{\rm sn}$
@@
where $G(\mathbf{k})\equiv[{\int\mathrm{d}^{3}{\bf x}~{}\bar{n}(\mathbf{x})w(\mathbf{x})e^{i\mathbf{k}\cdot\mathbf{x}}}]/[{[\int\mathrm{d}^{3}{\bf x}~{}\bar{n}^{2}(\mathbf{x})w^{2}(\mathbf{x})]^{1/2}}]$ is a window function [30, 31]. For $|{\bf k}|\gg 1/L$, where $L$ is a size of a survey window, $\langle\hat{P}(k)\rangle\simeq P(k)$ [e.g., 32], i.e. the estimator gives an unbiased estimate of the underlying power spectrum. Since the AM mocks and the single mass-threshold mocks have the same redshift dependence of halo distribution, $n_{\rm AM}(z)\propto n_{M_{\rm th}}(z)$, by design, the power spectra measured from these two mocks have the same effect of window convolution if we keep their $w(\mathbf{x})$ the same.
@@ 2.3.1 Linear limit of $P_{\rm AM}$
Similarly, in redshift space, we can express the local redshift-space power spectrum at position $x_{3}$ as $P_{\rm hh}^{\rm S}(\mathbf{k},M_{\rm th}(x_{3}))=[b_{1}^{2}(M_{\rm th}(x_{3}))+2f\mu^{2}b_{1}(M_{\rm th}(x_{3}))+f^{2}\mu^{4}]P^{\rm L}_{\rm mm}(k)$ in the linear regime, based on the Kaiser formula [13], where $f$ is the linear growth rate and $\mu$ is the cosine angle between the line of sight direction ($x_{3})$ and the wavenumber. Therefore the redshift-space power spectrum measured by the FKP estimator for the AM samples is
@@
$\overline{b_{1}}\equiv\frac{1}{\int\!\mathrm{d}x_{3}~{}n^{2}(x_{3})w^{2}(x_{3})}\int\!\mathrm{d}x_{3}~{}\bar{n}^{2}(x_{3})w^{2}(x_{3})b_{1}\!\left(M_{\rm th}\!(x_{3})\right).$ (2.12)

### v2

@@ 1 Introduction
[…]
@@
Therefore, this paper aims to investigate the impact of the selection effect on the redshift-space power spectrum. To achieve this, we consider a case where an assumed sample of galaxies follows a redshift distribution, denoted as $n(z)$ or $n(\chi)$, representing the number density of galaxies per unit redshift interval or per unit radial comoving distance interval. We do not consider the redshift evolution of galaxy clustering (structure formation) to focus on the impact of the redshift-dependent selection effect. Hence in our setting, the constant comoving number density, $n(\chi)\mathrm{d}V(\chi)={\rm const.}$, corresponds to a homogeneous sample or the absence of the selection effect. We first use the linear theory to show that the selection function generally causes a bias in the multipole moments of the redshift-space galaxy power spectrum, along the Feldman-Kaiser-Peacock estimator [21]. We then use halo catalogs from cosmological N-body simulations, AbacusSumitt [22, 23], to construct mock catalogs of galaxies (halos), by selecting halos based on a redshift-dependent mass threshold such that the resulting redshift distribution of the selected halos follows the target $n(z)$ for the SDSS BOSS LOWZ- or CMASS-like galaxies. We measure the multipole moments of the redshift-space power spectrum from the mock catalogs and then compare the moments with those obtained from halo catalogs of a single mass threshold, i.e. single tracers, in the same simulations. This comparison quantifies the impact of the selection effect on the redshift-space power spectrum.
@@ 2.1 N-body Simulations and Halo Catalogs
[…]
@@ 2.2.1 Mock catalogs with galaxy selection effect
To mimic actual galaxy surveys, we consider the LOWZ and CMASS galaxy samples of Sloan Digital Sky Survey Baryon Oscillation Spectroscopic Survey (BOSS) Data Release 11 (DR11) [^3] https://data.sdss.org/sas/dr11/boss/lss/; we consider LOWZ galaxies at $z>0.09$, and CMASS galaxies at $z>0.40$, respectively, as shown in Fig. 1. In addition to these cases we also consider a selection function of galaxies that follows a power-law selection of $n(z)$, given by $n(z)\propto z^{-\alpha}$, for the generality of our discussion. We investigate three cases, $\alpha=0.5,1.0,2.0$, and determine their normalizations so that the resulting averaged number density in the radial distance range of $0,1000]h^{-1}$Mpc, which roughly corresponds to the range of LOWZ, becomes the same as in the LOWZ case (see the right panel in Fig. 1). Throughout this paper we assume a distant observer approximation, and take the $z$-axis of each simulation box to be along the line-of-sight direction.
@@
Figure 1: The number density of LOWZ (left panel) and CMASS (middle panel) galaxies as a function of redshift, denoted as $n(z)$, taken from the SDSS DR11 catalog. Note that $n(z)$ is the number density per unit redshift interval. For illustrative purpose we also show the comoving $z$-coordinate on the bottom axis, computed using the inverse relation between redshift and the comoving radial distance at each redshift, $z=z^{-1}(\chi)$, based on the Planck 2018 cosmology. This $z$-coordinate can be compared to the size of AbacusSumitt simulation box ($2~h^{-1}$Gpc) used in this paper. The LOWZ galaxies are taken from the redshift range $z\simeq[0.09,0.5]$, while the CMASS galaxies are from the range $z\simeq[0.4,0.8]$. Note that we set the comoving $z$-coordinate at the lowest redshift to $\chi=0$ for each galaxy sample. We also consider a case of $n(z)$ following a power-law scaling given as $n(z)\propto z^{-\alpha}$ (right panel) to study the impact of galaxy selection effect as a general case.
@@
The LOWZ and CMASS galaxies are selected based on the color and magnitude cuts [19, 25, 26, 20]. It is very difficult and nearly impossible to exactly mimic the actual color-magnitude cut for the galaxy selection using the halo catalog, because it requires simulating galaxies while accounting for the complex physics of galaxy formation and evolution. In this paper, we employ the following simplified selection function as a working example. For a given $n(z)$ for the sample of galaxies considered, we select halos with masses above the mass threshold $M_{\rm th}(z)$ in each simulation realization so that the resulting number density matches the $n(z)$ in each redshift bin $\chi,\chi+\mathrm{d}\chi]$ (since the comoving radial distance is given by $\chi=\chi(z)$). We determine the mass threshold $M_{\rm th}(z)$ in each redshift bin from the average of the mass functions from 25 realizations of halo catalog, to minimize the sample variance effect. Since heavier halos tend to host more massive and therefore brighter galaxies [27], one could consider that this mass-threshold selection at each redshift mimics to some extent a flux limited sample of galaxies. In this paper, we refer to this method as abundance matching (AM) method, and denote the power spectrum measured from this catalog as $P_{\rm AM}(k)$. Fig. 2 shows the mass thresholds as a function of redshift, $M_{\rm th}(z)$, which we use to construct mock catalogs for LOWZ, CMASS and power-law samples, respectively.
@@
We construct each mock galaxy catalog in a rectangular-shape region of $2^{2}\times 1~(h^{-1}{\rm Gpc})^{3}$ volume, where we take the $z$-coordinate range of $0,1000]~h^{-1}$Mpc from each simulation box with a volume of $2^{3}~(h^{-1}{\rm Gpc})^{3}$, following Fig. 1. Thus we can generate two mock catalogs from each N-body simulation realization. We apply zero padding to the halo distribution to perform the Fast Fourier Transform in a $2^{3}~(h^{-1}{\rm Gpc})^{3}$ cubic box and then estimate the power spectrum using the Feldman-Kaiser-Peacock (FKP) estimator as described in Section 2.3 in detail. Since the mock catalog violates the periodic boundary condition, the estimated power spectrum is affected by the window convolution. For the following results, we use 50 realizations of the mock catalogs for each galaxy sample.
@@ 2.2.2 Mock catalogs with a single halo-mass threshold
To quantify the galaxy selection effect, we make a mock catalog without galaxy selection effect to be compared with the AM catalog. We create such a galaxy catalog using the following process:
@@
    In each simulation, from halos with masses above a given mass threshold, we randomly select halos in each redshift bin (or along the $z$-axis direction in our setting) so that the resulting redshift distribution of the number density follows $n(z)\propto\bar{n}_{\rm g}(z)$, where $\bar{n}_{\rm g}(z)$ is the redshift distribution for the LOWZ, CMASS or power-law galaxy sample, respectively, as shown in Fig. 1. When the number density of halos with masses above a given mass threshold is lower than the maximum number density of galaxies for each galaxy sample at a certain redshift, we down-sample the halos while maintaining the shape of $n(z)$. We estimate the power spectrum of the selected halos using the FKP estimator (see below), which retains the original power spectrum of the single mass threshold, and also incorporates the similar survey window effect as in the $P_{\rm AM}(k)$. From 25 realization, we obtain 50 halo catalogs (2 catalogs from each realization as described above) and their spectra.
@@
    Repeating the step (i), we find a certain mass threshold such that its (real-space) galaxy power spectrum obtained in step (i) matches the amplitude of the target spectrum $P_{\rm AM}(k)$ on linear scales in the range $k=[0.02,0.1]~h{\rm Mpc}^{-1}$. In this step we perform the $\chi^{2}$ fitting taking into account the sample variance error in each $k$ bin estimated from the variance of the 50 realizations. The halo catalog with the single mass threshold that minimizes the $\chi^{2}$ value is the catalog without a selection function, which we will use in the following analysis.
In the following we call this halo catalog as a single mass-threshold sample. We will compare the two power spectra measured from the AM sample and the single mass-threshold sample to quantify the impact of the galaxy selection effect.
Orange line in each panel of Fig. 3 shows $n(z)$ of a single mass-threshold sample, constructed based on the above method, for the LOWZ or CMASS galaxy sample. The dashed line in each panel shows the number density of halos corresponding to a certain mass threshold, which by definition has a constant amplitude in the $z$ direction because we are using the halo catalogs at the single redshift output ($z=0.5$). Then, we randomly select halos in each radial distance bin to satisfy $n(z)\propto n_{\rm LOWZ}(z)$ or $n_{\rm CMASS}(z)$, respectively. The proportional coefficient is determined by the ratio of the number density of halos of the single mass threshold (the dashed line in Fig. 3) to the peak number density of galaxies for each galaxy sample, $n_{\rm g}(z_{\rm peak})$; the number density at $z_{\rm peak}=0.08$ or $z_{\rm peak}\simeq 0.5$ for the LOWZ or CMASS sample, respectively. For LOWZ sample, the single mass threshold is $M_{\rm th}=9.14\times 10^{12}h^{-1}{\rm M_{\odot}}$, and the proportional factor in $n(z)\propto n_{\rm g}(z)$ is about $1/4$. For CMASS sample, the single mass-threshold is $M_{\rm th}=1.39\times 10^{12}h^{-1}{\rm M_{\odot}}$, and the proportional factor is about $1/2$. The power spectrum measured from this single mass-threshold sample includes the survey window effect similarly to that for the AM galaxy sample (see Section 2.3). Note that the shot noise contamination estimated by the FKP estimator is different between the AM and single mass-threshold samples, and we subtract it from the measured power spectrum.
@@
Figure 3: A figure illustrating the method for incorporating the window effect on the single-mass threshold sample, used to quantify the impact of selection effects in comparison to the results of the AM catalog. In each panel, the dashed line shows the number density which corresponds to the mass threshold in the legend. For the single mass-threshold sample, we randomly select halos above this single mass threshold in each redshift bin such that the redshift dependence of the number density is proportional to that for the LOWZ, CMASS or power-law sample in Fig. 1 (here we show results only the for the LOWZ and CMASS samples) (see Section 2.2.2 for details on the single mass-threshold sample). This single mass-threshold hold sample has the same window function as that for the AM sample.
@@ 2.3 FKP estimator
where $\bar{n}(\mathbf{x})$ is given by the redshift distribution $\bar{n}_{\rm g}(z)$ (Fig. 1) (our mocks do not have any variations in the mean number density in the $x,y$-directions), and we adopt $P_{0}=10^{4}~h^{-3}\rm{Mpc}^{3}$ for the AM sample. For the single mass-threshold sample, we set $P_{0}=10^{4}/\beta~h^{-3}\rm{Mpc}^{3}$ in Eq. (2.2), where the $\beta$ is the proportional factor of $\bar{n}_{M_{\rm th}}(z)\propto\bar{n}_{\rm g}(z)$ in Section 2.2.2. This scaling by $\beta$ is necessary to keep the window function (see Eq. 2.5) the same between AM sample and single mass-threshold sample. We generate the number density field of random particles $n_{s}(\mathbf{x})$ by repeating a Bernoulli process whose probability of success has the redshift dependence proportional to $\bar{n}_{g}(z)$, while the particles are randomly distributed in the $x$, $y$ directions.
@@
$\braket{\hat{P}(\mathbf{k})}$ $=\braket{|F(\mathbf{k})|{}^{2}}-P_{\rm sn}$
@@
where $G(\mathbf{k})\equiv[{\int\mathrm{d}^{3}{\bf x}~\bar{n}(\mathbf{x})w(\mathbf{x})e^{i\mathbf{k}\cdot\mathbf{x}}}]/[{[\int\mathrm{d}^{3}{\bf x}~\bar{n}^{2}(\mathbf{x})w^{2}(\mathbf{x})]^{1/2}}]$ is a window function [30, 31]. For $|{\bf k}|\gg 1/L$, where $L$ is a size of a survey window, $\langle\hat{P}(k)\rangle\simeq P(k)$ [e.g., 32], i.e. the estimator gives an unbiased estimate of the underlying power spectrum. Since the AM mocks and the single mass-threshold mocks have the same redshift dependence of halo distribution, $n_{\rm AM}(z)\propto n_{M_{\rm th}}(z)$, by design, the power spectra measured from these two mocks have the same effect of window convolution as long as we keep their $w(\mathbf{x})$ the same.
@@ 2.3.1 Linear limit of $P_{\rm AM}$
Similarly, in redshift space, we can express the local redshift-space power spectrum at position $x_{3}$ as $P_{\rm hh}^{\rm S}(\mathbf{k},M_{\rm th}(x_{3}))=[b_{1}^{2}(M_{\rm th}(x_{3}))+2f\mu^{2}b_{1}(M_{\rm th}(x_{3}))+f^{2}\mu^{4}]P^{\rm L}_{\rm mm}(k)$ in the linear regime, based on the Kaiser formula [13], where $f$ is the logarithmic growth rate and $\mu$ is the cosine angle between the line of sight direction ($x_{3})$ and the wavenumber. Therefore the redshift-space power spectrum measured by the FKP estimator for the AM samples is
@@
$\overline{b_{1}}\equiv\frac{1}{\int\!\mathrm{d}x_{3}~\bar{n}^{2}(x_{3})w^{2}(x_{3})}\int\!\mathrm{d}x_{3}~\bar{n}^{2}(x_{3})w^{2}(x_{3})b_{1}\!\left(M_{\rm th}\!(x_{3})\right).$ (2.12)